Evaluate the function for the given value of x:
1.) f(x) = 80-3x; f(5)
2.) f(x) = x(to the second power) + 4x -7; f(-1)
3.) f(x) = 3|x-4| + 2; f(2)

Answer by jennywecas(28)

1.) f(x) = 80-3x; f(5)
y=80-3(5)
y=80-15
y=65
2.) f(x) = x(to the second power) + 4x -7; f(-1)
y=-1*-1+4(-1)-7
y=1-4-7
y=-10
3.) f(x) = 3|x-4| + 2; f(2)
y=3|2-4|
y=3*2
y=6

btw if you need anymore help message me at jennywecas@yahoo.com